# A Scoping Review of Natural Language Processing in Addressing Medically Inaccurate Information: Errors, Misinformation, and Hallucination

### v1

@@ 6.1 Hallucination detection
In medical NLP, hallucination detection are critical for ensuring model reliability, particularly in tasks like question answering, summarization, and other text generation tasks. Hallucination detection focuses more on human evaluation of model-generated content rather than inaccuracies in existing texts. While error or misinformation detection generally uses classification metrics such as accuracy, precision, and recall, hallucination detection adds criteria like factual accuracy, coherence, faithfulness and hallucination rates.
@@
In other text generation tasks, hallucination detection focuses on ensuring the reliability of AI-generated academic and clinical content. Hua et al. [124] studied the accuracy of GPT-3.5 and GPT-4 when generating scientific abstracts and references in ophthalmology. Using modified DISCERN criteria (including helpfulness, truthfulness, and harmlessness), they evaluated quality and calculated hallucination rates by verifying AI-generated references. They found high hallucination rates in citations, suggesting the need for caution when using AI-generated academic content without verification. Liu et al. [131] introduced BenchHealth, a benchmark for testing LLM hallucination rates in healthcare across reasoning, generation, and understanding tasks. BenchHealth combined common metrics (e.g., accuracy, ROUGE-L, BLEU) with more specific measures of faithfulness, comprehensiveness, and robustness. This benchmark was used to evaluate general-purpose models like GPT-4 and fine-tuned medical LLMs like MedAlpaca. Results indicated that medical models often provided more faithful responses, while general-purpose models like GPT-4 offered more detailed answers with a higher risk of hallucination. Yang et al. [135] developed Scorpius, a conditional text generation system designed to test the vulnerability of medical knowledge graphs (KG) to hallucinated content. Scorpius generated malicious biomedical abstracts by using prompts for specific drug-disease pairs. These abstracts significantly increased drug relevance rankings within the KG. For example, 71.3 of drugs improved their rankings from the top 1,000 to the top ten positions. The quality of the Scorpius-generated abstracts was evaluated using metrics like perplexity, which showed better fluency and scientific consistency compared to ChatGPT. This study highlighted the potential risks posed by undetected hallucinated medical knowledge.
@@ 6.2 Hallucination mitigation
An intriguing perspective is presented by Qin et al. [132], who explored “patient hallucination” in doctor-patient dialogues. These hallucinations were defined as discrepancies between the symptoms expressed by the patient and their actual health conditions. They often arised from patients’ lack of medical knowledge, anxiety, or miscommunication, which potentially led to inaccurate or contradictory information during consultations. To tackle this issue, Qin et al. introduced MedPH, a medical dialogue generation framework that integrated both hallucination detection and mitigation. The detection module employed graph entropy analysis on a dynamic dialogue entity graph to identify three types of patient hallucination: isolated entities, denial of critical entities, and self-contradictions. For mitigation, MedPH generated clarifying questions informed by the hallucination-related context, guiding patients to articulate their conditions more accurately. Experimental results on medical dialogue datasets demonstrated that MedPH outperformed baseline models in both entity prediction and response generation tasks, significantly reducing hallucination rates while maintaining response quality. This approach highlights the importance of addressing patient-provided inaccuracies to enhance the reliability of medical dialogue systems.
@@ 7.2 Misinformation
Definitions. The distinction between unintentional misinformation and disinformation is often overlooked in existing studies. Without a clear understanding of the intent behind the content, it becomes difficult to assess the potential harm of the content and fully grasp its impact on public health.
@@
Future directions. Future research should focus on developing comprehensive, multimodal datasets that include diverse sources and formats. Improving context independence of claims in fact-checking tasks will enhance adaptability. NLP models should prioritize understanding the varying impact of misinformation on different subpopulations, focusing on vulnerable groups. Evaluating and implementing effective strategies and policies to prevent and mitigate health misinformation will advance both model performance and public health outcomes [142].

### v2

@@ 6.1 Hallucination detection
In medical NLP, hallucination detection is critical for ensuring model reliability, particularly in tasks like question answering, summarization, and other text generation tasks. Hallucination detection focuses more on human evaluation of model-generated content rather than inaccuracies in existing texts. While error or misinformation detection generally uses classification metrics such as accuracy, precision, and recall, hallucination detection adds criteria like factual accuracy, coherence, faithfulness and hallucination rates.
@@
In other text generation tasks, hallucination detection focuses on ensuring the reliability of AI-generated academic and clinical content. Hua et al. [126] studied the accuracy of GPT-3.5 and GPT-4 when generating scientific abstracts and references in ophthalmology. Using modified DISCERN criteria (including helpfulness, truthfulness, and harmlessness), they evaluated quality and calculated hallucination rates by verifying AI-generated references. They found high hallucination rates in citations, suggesting the need for caution when using AI-generated academic content without verification. Liu et al. [133] introduced BenchHealth, a benchmark for testing LLM hallucination rates in healthcare across reasoning, generation, and understanding tasks. BenchHealth combined common metrics (e.g., accuracy, ROUGE-L, BLEU) with more specific measures of faithfulness, comprehensiveness, and robustness. This benchmark was used to evaluate general-purpose models like GPT-4 and fine-tuned medical LLMs like MedAlpaca. Results indicated that medical models often provided more faithful responses, while general-purpose models like GPT-4 offered more detailed answers with a higher risk of hallucination. Yang et al. [137] developed Scorpius, a conditional text generation system that generates plausible yet hallucinated biomedical abstracts to test the vulnerability of medical knowledge graphs (KGs). Scorpius is built on an instruction-tuned LLM and guided by a scoring mechanism that refines synthetic abstracts to maximize their impact on KG rankings. By conditioning on specific drug-disease pairs, Scorpius generated abstracts that significantly elevated the relevance of target drugs - 71.3 of drugs improved their rankings from the top 1,000 to the top ten positions. The quality of the Scorpius-generated abstracts was evaluated using metrics like perplexity, which showed better fluency and scientific consistency compared to ChatGPT. This study highlighted the potential risks posed by undetected hallucinated medical knowledge.
@@ 6.2 Hallucination mitigation
An intriguing perspective is presented by Qin et al. [134], who explored “patient hallucination” in doctor-patient dialogues. These hallucinations were defined as discrepancies between the symptoms expressed by the patient and their actual health conditions. They often arose from patients’ lack of medical knowledge, anxiety, or miscommunication, which potentially led to inaccurate or contradictory information during consultations. To tackle this issue, Qin et al. introduced MedPH, a medical dialogue generation framework that integrated both hallucination detection and mitigation. The detection module employed graph entropy analysis on a dynamic dialogue entity graph to identify three types of patient hallucination: isolated entities, denial of critical entities, and self-contradictions. For mitigation, MedPH generated clarifying questions informed by the hallucination-related context, guiding patients to articulate their conditions more accurately. Experimental results on medical dialogue datasets demonstrated that MedPH outperformed baseline models in both entity prediction and response generation tasks, significantly reducing hallucination rates while maintaining response quality. This approach highlights the importance of addressing patient-provided inaccuracies to enhance the reliability of medical dialogue systems.
@@ 7.2 Misinformation
Definitions. The distinction between unintentional misinformation and disinformation is often overlooked in existing NLP studies. Without a clear understanding of the intent behind the content, it becomes difficult to assess the potential harm of the content and fully grasp its impact on public health. In clinical practice, failing to recognize intent can misguide interventions. A patient who misinterprets vaccine guidance due to low health literacy might benefit from personalized education [143], while combating anti-vaccine disinformation requires institutional efforts like public rebuttals or platform moderation [144]. At a societal level, disinformation can exacerbate health inequities by targeting vulnerable populations. For example, during the COVID-19 pandemic, disinformation campaigns disproportionately affected minority communities by spreading false claims about vaccine safety [145]. Without intent-aware models, NLP systems risk treating all inaccuracies as equal, which can lead to blunt mitigation strategies that fail to address the root cause of the misinformation.
@@
Future directions. Future research should focus on developing comprehensive, multimodal datasets that include diverse sources and formats. Improving context independence of claims in fact-checking tasks will enhance adaptability. NLP models should prioritize understanding the varying impact of misinformation on different subpopulations, focusing on vulnerable groups. To better differentiate unintentional misinformation from disinformation, future efforts should also prioritize the creation of benchmarking datasets annotated for intent (e.g., commercial, political, or sensational motives), which can guide the development of intent-aware classification models. Evaluating and implementing effective strategies and policies to prevent and mitigate health misinformation will advance both model performance and public health outcomes [147].